A MAS HAS 5 MORE ONE-DOLLAR BILLS THAN FIVE DOLLAR-BILLS, AND ONE MORE FIVE-DOLLAR BILL THAN TEN-DOLLAR BILLS. ALTOGETHER HE HAS $139 HOW MANY ONE-DOLLAR BILLS DOES HE HAVE?
I DONT THINK I AM SETTING UP THIS PROBLEM RIGHT.
THANKS

Answer by stanbon(75887)

A MAN HAS 5 MORE ONE-DOLLAR BILLS THAN FIVE DOLLAR-BILLS, AND ONE MORE FIVE-DOLLAR BILL THAN TEN-DOLLAR BILLS. ALTOGETHER HE HAS $139 HOW MANY ONE-DOLLAR BILLS DOES HE HAVE?
--------------------------
Let # of 5 dollar bills be "x".
Then # of 1 dollar bills is "x+5".
And # of 10 dollar bills "x-1".
-----------------------------------
Value Equation:
5x + 1(x+5) + 10(x-1) = 139
5x + x + 5 + 10x - 10 = 139
16x = 144
x = 9 (# of 5 dollars bills)
# of 1 dollar bills = x+5 = 14
# of 10 dollar bills = x-1 = 8
=================================
Cheers,
Stan H.